OpenStudy (anonymous):

If a,b,c, and f are four nonzero numbers, then all of the following proportions are equivalent EXCEPT a/f=b/c, f/c=b/a, c/a=f/b, a/c =b/f, and af/bc=1/1

OpenStudy (anonymous):

so first we start with the information that we can use: since we know that: \[af/bc = 1/1\] then: \[af= bc\] now from the above statement we can make the following expressions: first, if we divide both sides by c and both sides again by f, we get: \[a/c= b/f\] Second, if we divide both sides by a and both sides again by b, we get: \[f/b= c/a\] Third, if we divide both sides by a and both sides again by c, we get: \[f/c= b/a\] Now, the only form we will not be able to get is: \[a/f= b/c\] just because if we cross multiply, we will get: \[ac=fb\] which does not agree with the first statement we made at the beginning, that is: IF: \[af/bc = 1/1\] then: \[af= bc\] Do you get it?

OpenStudy (anonymous):

yeah, i get it. thanks :D